The ages of Ann and Bill now add up to 91. Ann is three times as old as Bill was at the time when Ann was twice as old as Bill is now. How old are Ann and Bill?
Answer by ankor@dixie-net.com(12695)

The ages of Ann and Bill now add up to 91. Ann is three times as old as Bill was at the time when Ann was twice as old as Bill is now. How old are Ann and Bill?
:
Write an equation for each statement:
"The ages of Ann and Bill now add up to 91."
a + b = 91
:
"Ann is three times as old as Bill was at the time when Ann was twice as old as Bill is now."
a = 3(2b)
a = 6b
:
How old are Ann and Bill?
Substitute 6b for a in the sum equation
6b + b = 91
b = 91/7
b = 13 yrs is Bill's age
then
91 - 13 = 78 yrs is Ann's age
:
Check solution in the statement:
" Ann is three times as old as Bill was at the time when Ann was twice as old as Bill is now. "
78 = 3(2*13)
78 = 3(26)